OpenStudy (anonymous):

Simplify the complex rational expression using the LCD. ((9/7)-(1/77))/((6/35)+(1/7))

OpenStudy (lgbasallote):

6/35 + 1/7 = 6 + 5/ 35 = 11/35 9/7 - 1/77/11/35 1/77/11/35 = 1/77 x 35/11 = 1/11 x 5/11 = 5/121 9/7 - 5 / 121 1089 -35/847 =1054/847 1 207/847

OpenStudy (anonymous):

Thank you so much lgbasallote!! :)

OpenStudy (anonymous):

question, so the answer would be 1207/847? or 1054/847?

OpenStudy (lgbasallote):

both are equal...if they want a mixed give them the mixed...if they want the improper give them the improper

OpenStudy (anonymous):

ooh alright. thank you so much!